5. The average number of units of an item produced by a machine in 1 minute is 10. Find the probability that:
i. Machine produces 5 units
ii. Machine produces at least 2 units

Answer by Boreal(15235)

Poisson with parameter 10. counts, proportional to time, can be theoretically infinite
e^(-10)^10^5/5!=0.0378
at least 2 units, look at 1 unit and 0, get those probabilities and subtract from 1.
p(0)=e^(-10)=0.000045
p(1)=e^(-10)10*1/1!=0.00045
These two add to 0.0005
probability at least 2 is 0.9995